# Arnold-Thom conjecture for the arrival time of surfaces

### v1

@@ 2.2. Neck singularities
Let $M_{t}$ be a mean convex mean curvature flow, $u$ be its arrival time function, and $\gamma(t)$ be a gradient flow line of $u$ converging to a singular point $p.$ Suppose $\gamma$ is parametrized by the time of the flow. If there are sequences $t_{i}\to 0$ and $\lambda_{i}\to\infty$ such that the sequence of rescaled flows $M^{t_{i}}_{t}:=M^{\left(\gamma(t_{i}),t_{i},\lambda_{i}\right)}_{t}$ converges to an ancient oval or a shrinking sphere, then the limit point of the flow line is a spherical singularity.
@@
We can now state the following convergence result for a sequence of flows obtained by rescaling at a sequence of points along a gradient flow line of an arrival time function. It follows directly from Lemma 2.6, [HK1, Theorem 1.12], [CHHW, Theorem 9.1] (cf. [SS, Proposition 2.3]), and the fact that $H\left(\gamma(t),t\right)\to\infty$ if $\gamma(t)$ converges to a critical point of $u.$ Here, we use $H(x,t)$ to get a sequence of rescaled flows instead of $1/R(x,t)$ used in [CHHW] since they are equivalent in the setting of mean convex flows.
@@
Let $M_{t}$ be a mean convex mean curvature flow, $u$ be its arrival time function, and $\gamma(t)$ be a gradient flow line of $u$ converging to a singular point $p.$ If $p$ is a neck singularity, then any limit flow along $\gamma$ is a shrinking cylinder or a translating bowl with its axis parallel to that of the tangent flow. To be more specific, if $\gamma$ is parametrized by time, then given any $t_{i}\to 0,$ after passing to a subsequence, the sequence of flows $M^{t_{i}}_{t}:=M^{\left(\gamma(t_{i}),t_{i},\lambda_{i}\right)}_{t},$ where $\lambda_{i}=H\left(\gamma(t_{i}),t_{i}\right),$ converges to a non-collapsed ancient solution, which is either a shrinking cylinder or a translating bowl soliton with its axis parallel to that of the cylindrical tangent flow.
A subtle point is that in Proposition 2.8, we assume the convergence of $\gamma,$ which is stronger than the assumption of just having a limit point. This will be fine since later we will first prove the Łojasiewicz theorem for such a flow line, and hence having a limit point implies that the flow line converges.
@@
A general version of the backward improvement was done by Gianniotis-Haslhofer in [GH, Proposition 4.2], where they dealt with all ancient non-collapsed mean curvature flows. Here, we only need the result when the model is a bowl soliton or a cylinder, so we can get a stronger improvement statement.
We end this section by recording a convergence result of flow lines when looking at limit flows along $\gamma.$ This lemma follows from a general statement in the appendix (Lemma A.1).
@@ 3.1. Uniqueness of limit points
First, we prove the Łojasiewicz theorem for $\gamma$. This is based on the structure theorems of cylindrical singularities by Colding-Minicozzi [CM162] (cf. [G24]) and some classical results about limit points of $C^{1}$ curves. [^5] In the theory of ordinary differential equations, the set of limit points of a gradient flow line is an interesting object. One can, for instance, see [P01, Section 3.2], based on which the arguments about the connectedness in the proof are structured.
@@
Let $K$ be the set of limit points of $\gamma.$ That is,
$K:=\left\{x\in\Omega:\gamma\left([-T,0)\right)\cap B_{\delta}(x)\setminus\left\{x\right\}\neq\emptyset\text{ for all }\delta>0\right\}.$
@@
We prove the claim first. Suppose $x\in\overline{K}.$ It suffices to show that given any $\delta>0,$ we have
(3.3) $\gamma\left([-T,0)\right)\cap B_{\delta}\neq\emptyset.$
First, since $x\in\overline{K},$ we can find $y\in K\cap B_{\delta/2}(x)\setminus\left\{x\right\}.$ Let $d:=|x-y|>0.$ Because $y\in K,$ we can find $t\in[-T,0)$ such that $\gamma(t)\in B_{d/2}(y)\setminus\left\{y\right\}.$ Then we have $\gamma(t)\in B_{d/2}(y)\subseteq B_{\delta}(x).$ Moreover, $|\gamma(t)-x|\geq\delta/2-d/2>0,$ so $\gamma(t)\neq x.$ Thus, $\gamma(t)\in B_{\delta}(x)\setminus\left\{x\right\},$ and hence we get (3.3). This proves that $K$ is closed. Since $\Omega$ is bounded and $K\subseteq\Omega,$ $K$ is compact.
@@
Now, we prove the conclusion by contradiction. Suppose $|K|>1$, i.e., limit points are not unique. Let $p\in K$ be a neck singularity. The structure theorem in [CM162] then implies that $K\cap\overline{B}_{\delta}(p)$ consists of only spherical and neck singularities and is contained in a $C^{1}$ curve when $\delta>0$ is small. Since $K$ is a compact connected set, after shrinking $\delta$ if necessary, $K\cap\overline{B}_{\delta}(p)$ is also compact and connected and hence itself forms a $C^{1}$ curve (instead of being a proper subset of it) if $\left|K\cap\overline{B}_{\delta}(p)\right|>1.$ Since the arrival time function is $C^{2}$ near a spherical singularity by Huisken’s result [H93], every point in $K\cap\overline{B}_{\delta}(p)$ must be a neck singularity. If this $C^{1}$ curve had an interior point, the arrival time function would be $C^{2}$ at this point by Theorem 2.4, and the Łojasiewicz inequality [CM19, Theorem 2.1] implies the uniqueness of limit points, that is, $|K|=1,$ a contradiction. Hence, $K\cap\overline{B}_{\delta}(p)$ consists of one point, so $K$ only contains isolated points. Since $K$ is connected, the only possibility is $|K|=1.$ This again contradicts the assumption that limit points are not unique. Therefore, the conclusion follows. ∎
@@ 3.2. Uniqueness of limit flows
We start to look at the two types of limit flows along $\gamma.$ By Proposition 2.8, they are either shrinking cylinders or translating bowl solitons. First, we show a backward improvement result for cylindricality.
@@
If we were not able to find such a $\tau<0$, then we could find $t_{i}\to 0$ such that $\left(\gamma(t_{i}),t_{i}\right)$ is $\delta$-cylindrical but either $(\gamma(t),t)$ is not $\varepsilon$-cylindrical when $t\in\left[t_{i}-H(t_{i})^{-2}\mathcal{T},t_{i}\right]$ or $(\gamma(t),t)$ is not $\delta$-cylindrical at $t=t_{i}-H(t_{i})^{-2}\mathcal{T}.$ By Proposition 2.8, we can pass to a subsequence such that
@@
where $M_{t}^{\infty}$ is a translating bowl soliton or a shrinking cylinder. By the assumption, $(0,0)$ is $2\delta$-cylindrical in $M_{t}^{\infty}$. By Lemma 2.9, there is a gradient flow line $\gamma^{\infty}(t)$ of $M_{t}^{a}$ for $t\in(-\infty,0]$ with $\gamma^{\infty}(0)=0,$ and Lemma 2.11 implies that the rescaled flow line
@@
By the claim (3.5), $(\gamma(t),t)$ is $\varepsilon$-cylindrical when $t\in[\overline{t}_{i+1},\overline{t}_{i}]$ and $(\gamma(\overline{t}_{i}),\overline{t}_{i})$ is $\delta$-cylindrical. Therefore, $\partial_{t}H=\Delta H+|A|^{2}H>0$ and consequently $H(t)=H(\gamma(t),t)$ is strictly increasing on $[\overline{t}_{i},\overline{t}_{1}].$ This means that $\overline{t}_{j}<\overline{t}_{1}-(j-1)H\left(\overline{t}\right)^{-2}\mathcal{T}.$ Therefore, the process must stop at $\overline{t}_{\ell}$ for some finite $\ell$. Now, we use the claim again to conclude that $(\gamma(t),t)$ is $\varepsilon$-cylindrical when $t\in\left[\overline{t}_{l}-H\left(\overline{t}_{l}\right)^{-2}\mathcal{T},\overline{t}\right]$. Since $\overline{t}_{l}-H\left(\overline{t}_{l}\right)^{-2}\mathcal{T}\leq\tau,$ the assertion of the lemma follows. ∎
@@
Next, we fix an arbitrary $\varepsilon>0$ and take the constants $\delta$ and $\tau$ from Lemma 3.4. Since 00 is not the tip, by Lemma 2.10, there exists $\mathcal{T}<\infty$ such that $(\overline{\gamma}(t),t)$ is $\frac{\delta}{2}$-cylindrical when $t\leq-\mathcal{T}.$ By Lemma 2.11, $\gamma^{t_{i}}(t)$ converges uniformly to $\overline{\gamma}(t)$ when $t\in[-2\mathcal{T},0].$ Rescaling back implies that $\left(\gamma(t_{i}-H(t_{i})^{-2}\mathcal{T}),t_{i}-H(t_{i})^{-2}\mathcal{T}\right)$ is $\delta$-cylindrical for all sufficiently large $i$. By Lemma 3.4, $(\gamma(t),t)$ is $\varepsilon$-cylindrical when $t\in[\tau,t_{i}-H(t_{i})^{-2}\mathcal{T}).$ Since $t_{i}$ and $H(t_{i})^{-1}$ both converge to $0,$ this implies that $(\gamma(t),t)$ is $\varepsilon$-cylindrical when $t\in[\tau,0).$ In particular, this means that $M^{\infty}_{t}$ is $2\varepsilon$-cylindrical at $(0,0).$ Since $\varepsilon$ is arbitrary, we conclude that $M^{\infty}_{t}$ is a shrinking cylinder.
@@ 3.3. Uniqueness of limit tangents
Based on Theorem 3.8 and Theorem 3.10, we may call a flow line cylinder-type or bowl-type depending on the limit flows along it. Based on the types of limit flows along a gradient curve limiting to a neck singularity, there are two possibilities. That is, it is either cylinder-type or bowl-type. For the former case, we will reduce it to the work of Colding-Minicozzi [CM18, CM19]; for the latter case, we will deal with it directly. Using the uniqueness of limit flows, we can prove the main theorem.
Theorem 3.11 (Arnold-Thom for neck singularities).
This immediately implies Theorem 1.3 and the case of two-convex flows claimed in Section 1.
Proof.
The uniqueness of limit points is addressed in Theorem 3.1. For the uniqueness of limit tangents, we only need to deal with the case when the limit point is a neck singularity since the arrival time function is $C^{2}$ at spherical points. We may assume the limit point is the origin and we can do a time translation so that it happens at time $t=0$. In particular, $\lim\limits_{t\rightarrow 0}\gamma(t)=0$.
There are two cases based on the types of limit flows along $\gamma.$
Case 1: A limit flow along $\gamma$ is a translating bowl soliton.
Case 2: A limit flow along $\gamma$ is a shrinking cylinder.
We let $e_{0}\in\left\{0\right\}\times\mathbb{R}\subseteq\mathbb{R}^{n}\times\mathbb{R}$ be the unit vector parallel to the axis of the tangent flow. By Theorem 3.8, given any $t_{i}\to 0,$ after passing to a subsequence, the limit flow is a shrinking cylinder parallel to $e_{0}$.
Let $\Pi_{\rm axis}$ be the orthogonal projection onto $\left\{0\right\}\times\mathbb{R}$ and $\Pi:={\rm Id}_{\mathbb{R}^{n+1}}-\Pi_{\rm axis}$, where $\mathbb{R}$ is the axis of the tangent flow, which is parallel to $e_{0}$. Therefore, $\gamma^{\prime}=\textbf{H}$ behaves similarly to the shrinking cylinder near the limit point. More precisely, there exists $\tau<0$ such that for all $t\in[\tau,0)$ we have
@@
(3) $\left|\Pi_{\rm axis}\circ\gamma^{\prime}(t)\right|=|\langle\gamma^{\prime},e_{0}\rangle(t)|=o\left(|\gamma^{\prime}|\right).$
@@
Combining (3.13) and (3.14), we see that for any $d>0,$ $\gamma(t)$ is in the region
$C_{d}:=\left\{x\in\Omega:|\Pi_{\rm axis}(x)|\leq d|\Pi(x)\right\}$
when $t\in[\tau,0).$ For such $d>0,$ by [CM18, Theorem 1.3], there exists $\delta>0$ such that $u$ is $C^{2}$ and has the unique critical point 00 in $C_{d}\cap B_{\delta}.$ The proofs of [CM19, Theorems 2.1, 2.2, 3.3] then go through in the region $C_{d}\cap B_{\delta}$ since the key property is the $C^{2}$-regularity of $u.$ The summability estimate [CM19, Equation (3.17)] was proven in [CM19, Sections 4, 5, and 6] for any rescaled mean curvature flow converging to a generalized cylinder. Thus, it is true in our case. This completes the proof. ∎

### v2

@@ 2.2. Neck singularities
Let $M_{t}$ be a mean convex mean curvature flow, $u$ be its arrival time function, and $\gamma(t)$ be a gradient flow line of $u$ converging to a singular point $p$ at time $0,$ that is, $u(p)=0.$ Suppose $\gamma$ is parametrized by the time of the flow. If there are sequences $t_{i}\to 0$ and $\lambda_{i}\to\infty$ such that the sequence of rescaled flows $M^{t_{i}}_{t}:=M^{\left(\gamma(t_{i}),t_{i},\lambda_{i}\right)}_{t}$ converges to an ancient oval or a shrinking sphere, then the limit point of the flow line is a spherical singularity.
@@
We can now state the following convergence result for a sequence of flows obtained by rescaling at a sequence of points along a gradient flow line of an arrival time function.
@@
Let $M_{t}$ be a mean convex mean curvature flow, $u$ be its arrival time function, and $\gamma(t)$ be a gradient flow line of $u$ converging to a singular point $p$ with $u(p)=0.$ If $p$ is a neck singularity, then any limit flow along $\gamma$ is a shrinking cylinder or a translating bowl with its axis parallel to that of the tangent flow. To be more specific, if $\gamma$ is parametrized by time, then given any $t_{i}\to 0,$ after passing to a subsequence, the sequence of flows $M^{t_{i}}_{t}:=M^{\left(\gamma(t_{i}),t_{i},\lambda_{i}\right)}_{t},$ where $\lambda_{i}=H\left(\gamma(t_{i}),t_{i}\right),$ converges to a non-collapsed ancient solution, which is either a shrinking cylinder or a translating bowl soliton with its axis parallel to that of the cylindrical tangent flow.
Since this result is essential in the rest of the paper, we briefly outline how it follows from the previous results mentioned above. One can also see Appendix B for a complete proof.
Proof outline of Proposition 2.8.
The proposition follows directly from Lemma 2.6 and [CHHW, Proposition 9.1] (cf. [SS, Proposition 2.3]). Here, we use $H(x,t)$ to rescale the flow and get a blowup limit instead of using $1/R(x,t)$ in [CHHW] since they are equivalent in the setting of mean convex flows based on the curvature estimates [HK1, Theorem 1.8].
For the fact about parallel axes, it was first mentioned at the end of the proof outline [CHHW, Section 1.7]. A proof can be found in the proof of [SS, Proposition 2.3(iv)]. This improvement is based on the fact that in the definition of the cylindrical scale in [SS], the information of axes is kept.
The proof in [SS] essentially provides what we outline above. To be self-contained, a complete proof is given in Appendix B for interested readers. ∎
A subtle point in Proposition 2.8 is that we assume the convergence of $\gamma,$ which is stronger than the assumption of just having a limit point. This will not be an issue since later we will first prove the Łojasiewicz theorem for such a flow line (Theorem 3.1), and hence having a limit point implies that the flow line converges.
@@
A general version of the backward improvement was done by Gianniotis–Haslhofer in [GH, Proposition 4.2], where they dealt with all ancient non-collapsed mean curvature flows. Here, we only need the result when the model is a bowl soliton or a cylinder, so we can get a stronger improvement statement.
We end this section by recording a convergence result of flow lines when looking at limit flows along $\gamma.$ This lemma follows from a general statement in Appendix A (Lemma A.1).
@@ 3.1. Uniqueness of limit points
First, we prove the Łojasiewicz theorem for $\gamma$. This is based on the structure theorems of cylindrical singularities by Colding–Minicozzi [CM162] (cf. [G24]) and some classical results about limit points of $C^{1}$ curves. [^5] In the theory of ordinary differential equations, the set of limit points of a gradient flow line is an interesting object. One can, for instance, see [P01, Section 3.2], based on which the arguments about the connectedness in the proof are structured.
@@
We remark that a possible alternative approach to proving the uniqueness of limit points is to appeal to the uniqueness of limit flows, which will be proved in Theorems 3.7 and 3.9. See Remark 4.20. However, in contrast to the current version of Theorem 3.9, this alternative approach requires additional work to establish the uniqueness of the axes of bowl solitons, since the limit point may not be unique a priori. Here, we instead opt for a direct proof, as the structure theory of singular sets developed by Colding–Minicozzi [CM162] is particularly effective in the presence of spherical and neck singularities. This strategy also takes advantage of the special structure of one-dimensional curves and avoids difficulties arising from non-uniqueness of limit points, thereby simplifying the preparation in Section 3.2 for the proof of the Arnold–Thom conjecture.
@@
Let $K$ be the set of limit points of $\gamma(t)$ as $t\to 0.$ That is, [^6] One can also define $K$ by looking at the closure of the spacetime track of $\gamma$ intersected with the time-zero slice. With this definition, the closedness will follow directly.
$K:=\left\{x\in\Omega:\lim_{i\to\infty}\gamma(t_{i})=x\text{ for some }t_{i}\nearrow 0\right\}.$
@@
We prove the claim first. To show the compactness, it suffices to show that $K$ is closed since $K$ is a bounded region. The closedness then follows from a standard diagonal argument. In fact, if $x\in\overline{K},$ then we can find $x^{j}\in K$ with $x^{j}\to x$ as $j\to\infty.$ For each $j,$ we can find $t^{j}_{i}\in[-T,0)$ such that $\gamma(t^{j}_{i})\to x^{j}$ as $i\to\infty.$ We may assume for each $j,$ $t^{j}_{i}$ is increasing. Then for each $j,$ we take $i(j)\in\mathbb{N}$ such that
$d_{\rm euc}\left(\gamma\left(t^{j}_{i(j)}\right),x^{j}\right)\leq\frac{1}{j}.$
The triangle inequality then implies $\gamma\left(t^{j}_{i(j)}\right)\to x$ as $j\to\infty.$ This proves that $K=\overline{K}.$ Thus, $K$ is closed and hence compact.
@@
Now, we prove the conclusion by contradiction. Suppose #$\#K>1$, i.e., limit points are not unique. Let $p\in K$ be a neck singularity. The structure theorem in [CM162] then implies that $K\cap\overline{B}_{\delta}(p)$ consists of only spherical and neck singularities and is contained in a $C^{1}$ curve when $\delta>0$ is small. Since $K$ is a compact connected set, after shrinking $\delta$ if necessary, $K\cap\overline{B}_{\delta}(p)$ is also compact and connected and hence itself forms a $C^{1}$ curve (instead of being a proper subset of it) if #$\#\left(K\cap\overline{B}_{\delta}(p)\right)>1.$ Since the arrival time function is $C^{2}$ near a spherical singularity by Huisken’s result [H93], every point in $K\cap\overline{B}_{\delta}(p)$ must be a neck singularity. If this $C^{1}$ curve had an interior point, the arrival time function would be $C^{2}$ near this point by Theorem 2.4, and the Łojasiewicz inequality [CM19, Theorem 2.1] implies the uniqueness of limit points, that is, #$\#K=1,$ a contradiction. Hence, $K\cap\overline{B}_{\delta}(p)$ consists of one point, so $K$ only contains isolated points. Since $K$ is connected, the only possibility is #$\#K=1.$ This again contradicts the assumption that limit points are not unique. Therefore, the conclusion follows. ∎
@@ 3.2. Uniqueness of limit flows
We start to look at the two types of limit flows along $\gamma.$ By Proposition 2.8, they are either shrinking cylinders or translating bowl solitons. As mentioned, a mean convex mean curvature flow $M_{t}$ ($t\in[-T,\infty)$) is fixed. First, we show a backward improvement result for cylindricality.
@@
We remark that the constants $\delta$ and $\tau$ depend on $M_{t},$ and we can apply the lemma to other mean convex flows.
@@
If we were not able to find such a $\tau<0$, then we could find $t_{i}\to 0$ such that $\left(\gamma(t_{i}),t_{i}\right)$ is $\delta$-cylindrical but either $(\gamma(t),t)$ is not $\varepsilon$-cylindrical for some $t\in\left[t_{i}-H(t_{i})^{-2}\mathcal{T},t_{i}\right]$ or $(\gamma(t),t)$ is not $\delta$-cylindrical at $t=t_{i}-H(t_{i})^{-2}\mathcal{T}.$ By Proposition 2.8, we can pass to a subsequence such that
@@
where $M_{t}^{\infty}$ is a translating bowl soliton or a shrinking cylinder. By the assumption, $(0,0)$ is $2\delta$-cylindrical in $M_{t}^{\infty}$. By Lemma 2.9, there is a gradient flow line $\gamma^{\infty}(t)$ of $M_{t}^{\infty}$ for $t\in(-\infty,0]$ with $\gamma^{\infty}(0)=0,$ and Lemma 2.11 implies that the rescaled flow line
@@
By the claim (3.4), $(\gamma(t),t)$ is $\varepsilon$-cylindrical when $t\in\left[\overline{t}_{i+1},\overline{t}_{i}\right]$ and $(\gamma(\overline{t}_{i}),\overline{t}_{i})$ is $\delta$-cylindrical. Therefore,
$\partial_{t}H=\Delta H+|A|^{2}H\geq\left(\frac{1}{n-1}-C\varepsilon\right)H^{3}>0,$
noting that $\Delta H_{\rm cyl}=0$ and $|A_{\rm cyl}|^{2}H_{\rm cyl}=\frac{1}{n-1}H_{\rm cyl}^{3}$ on an exact cylinder $S^{n-1}\times\mathbb{R}$. Consequently, $H(t)=H(\gamma(t),t)$ is strictly increasing on $\left[\overline{t}_{i},\overline{t}_{1}\right].$ This means that $\overline{t}_{j}<\overline{t}_{1}-(j-1)H\left(\overline{t}\right)^{-2}\mathcal{T}.$ Therefore, the process must stop at $\overline{t}_{\ell}$ for some finite $\ell$. Now, we use the claim again to conclude that $(\gamma(t),t)$ is $\varepsilon$-cylindrical when $t\in\left[\overline{t}_{l}-H\left(\overline{t}_{l}\right)^{-2}\mathcal{T},\overline{t}\right]$. Since $\overline{t}_{l}-H\left(\overline{t}_{l}\right)^{-2}\mathcal{T}\leq\tau,$ the assertion of the lemma follows. ∎
@@
Next, we fix an arbitrary $\varepsilon>0$ and take the constants $\delta$ and $\tau$ from Lemma 3.3 applied to $M^{\infty}_{t}$. Since 00 is not the tip, by Lemma 2.10, there exists $\mathcal{T}<\infty$ such that $(\overline{\gamma}(t),t)$ is $\frac{\delta}{2}$-cylindrical when $t\leq-\mathcal{T}.$ By Lemma 2.11, $\gamma^{t_{i}}(t)$ converges uniformly to $\overline{\gamma}(t)$ when $t\in[-2\mathcal{T},0].$ Rescaling back implies that $\left(\gamma(t_{i}-H(t_{i})^{-2}\mathcal{T}),t_{i}-H(t_{i})^{-2}\mathcal{T}\right)$ is $\delta$-cylindrical for all sufficiently large $i$. By Lemma 3.3, $(\gamma(t),t)$ is $\varepsilon$-cylindrical when $t\in[\tau,t_{i}-H(t_{i})^{-2}\mathcal{T}).$ Since $t_{i}$ and $H(t_{i})^{-1}$ both converge to $0,$ this implies that $(\gamma(t),t)$ is $\varepsilon$-cylindrical when $t\in[\tau,0).$ In particular, this means that $M^{\infty}_{t}$ is $2\varepsilon$-cylindrical at $(0,0).$ Since $\varepsilon$ is arbitrary, we conclude that $M^{\infty}_{t}$ is a shrinking cylinder.
@@ 3.3. Uniqueness of limit tangents
Based on Theorem 3.7 and Theorem 3.9, we may call a flow line cylinder-type or bowl-type depending on the limit flows along it. Based on the types of limit flows along a gradient curve limiting to a neck singularity, there are two possibilities. That is, it is either bowl-type or cylinder-type. For the former case, we will deal with it directly; for the later case, we will reduce it to the work of Colding–Minicozzi [CM18, CM19]. Since we will work in a similar situation in Section 4.3, we single out this result for cylinder-type singularities for all types of generalized cylinders.
Lemma 3.10.
Let $M_{t}$ be a closed mean convex mean curvature flow in $\mathbb{R}^{n+1}$ and let $u$ be its arrival time function with $u(0)=0$. Suppose 00 is a cylindrical singularity whose tangent flow is a $k$-cylinder and suppose $\gamma$ is a gradient flow line for $u$ converging to $0$ If all limit flows along $\gamma$ are $k$-cylinders with axes parallel to that of the tangent flow, then the unit tangents of $\gamma$ converge.
This is a consequence of the geometry of shrinking cylinders and theories established by Colding–Minicozzi. We will apply Lemma 3.10 with $k=1$ in this section and with general $k$ in Section 4.3. We remark again that Lemma 3.10 already assumes the convergence of the flow line, which is fine since we always prove the Łojasiewicz theorem first before applying this lemma both here and in Section 4.3. In Section 4, we make use of similar techniques in the proof of Theorem 4.13, while different aspects of the evolution inequalities are explored. See Remark 4.21.
Proof of Lemma 3.10.
After a rotation, we may assume the axis of the tangent flow is $\left\{0\right\}\times\mathbb{R}^{k}$ where $0=0^{n-k+1}\in\mathbb{R}^{n-k+1}.$ Let $\Pi_{\rm axis}$ be the orthogonal projection onto $\left\{0\right\}\times\mathbb{R}^{k}$ and define its complement $\Pi:={\rm Id}_{\mathbb{R}^{n+1}}-\Pi_{\rm axis}.$
The assumption implies that $\gamma^{\prime}=\textbf{H}$ behaves similarly to that of the shrinking cylinder near the limit point. More precisely, there exists $\tau<0$ such that for all $t\in[\tau,0)$ we have
@@
(3) $\left|\Pi_{\rm axis}\circ\gamma^{\prime}(t)\right|=|\langle\gamma^{\prime},e_{0}\rangle(t)|=o\left(|\gamma^{\prime}|\right)$ as $t\to 0.$
@@
Combining (3.12) and (3.13), we see that there exists $d<\infty$ large enough such that
(3.14) $\gamma(t)\in C_{d}:=\left\{x\in\Omega:|\Pi_{\rm axis}(x)|\leq d|\Pi(x)|\right\}$
for $t\in[\tau,0).$
Now, we are going to apply the theories established by Colding–Minicozzi in [CM18, CM19]. First, by [CM18, Theorem 1.3], there exists $\delta>0$ such that $u$ is $C^{2}$ and has the unique critical point 00 in $C_{d}\cap B_{\delta}.$ The proofs of [CM19, Theorems 2.1, 2.2, 3.3] then go through in the region $C_{d}\cap B_{\delta}$ since the key property is the $C^{2}$-regularity of $u.$ To be more precise, the proof of [CM19, Theorem 2.1], which is essentially based on the fundamental theorem of calculus, implies that given any $\overline{\varepsilon}>0,$ we can find $\overline{\delta}\in(0,\delta)$ such that
$\left|\nabla u(q)+\frac{q-p}{n-k}\right|\leq\overline{\varepsilon}|q-p|$
for all $p,q\in C_{d}\cap B_{\overline{\delta}}.$ Since this is true for any $\overline{\varepsilon}>0,$ in particular, the inequality implies
(3.15) $\frac{|\nabla u|^{2}}{-u}(p)\to\frac{2}{n-k}$
as $p\in C_{d}\cap B_{\delta}$ tends to $0$ The relation (3.15) and the property of being $C^{2}$ are the only two ingredients in the proofs of [CM19, Theorems 2.2 and 3.3]. Thus, [CM19, Theorem 3.3] is valid in our case by (3.14) and [CM18, Theorem 1.3].
With [CM19, Theorem 3.3], we know that if the summability estimate [CM19, Equation (3.17)] is true for the rescaled mean curvature flow obtained by rescaling with respect to a cylindrical singularity $(0,0),$ then Theorem 3.16 follows. Thus, the summability estimate [CM19, Equation (3.17)] is the only missing piece. However, in [CM19, Sections 4, 5, and 6], Colding–Minicozzi proved [CM19, Equation (3.17)] for any rescaled mean curvature flow converging to a generalized cylinder without additional conditions (like $C^{2}$-regularity of $u$). In particular, the estimate [CM19, Equation (3.17)] is true for the rescaled mean curvature flow obtained with respect to the cylindrical singularity $(0,0)$ in our case. This proves Lemma 3.10. ∎
Now, we use the uniqueness of limit flows and Lemma 3.10 to prove the main theorem of this section. It immediately implies Theorem 1.3 and the case of two-convex flows claimed in Section 1.1 since the arrival time function is $C^{2}$ at spherical points and hence Colding–Minicozzi’s work already covers this case.
Theorem 3.16 (Arnold–Thom for neck singularities).
Proof.
The uniqueness of limit points is addressed in Theorem 3.1. For the uniqueness of limit tangents, we may assume the limit point is the origin and that it happens at time $t=0$. In particular, $\lim\limits_{t\rightarrow 0}\gamma(t)=0$.
There are two cases based on the types of limit flows along $\gamma.$
Case 1: A limit flow along $\gamma$ is a translating bowl soliton.
Case 2: A limit flow along $\gamma$ is a shrinking cylinder.
By Theorem 3.7, given any $t_{i}\to 0,$ after passing to a subsequence, the limit flow is a shrinking cylinder parallel to the axis of the neck tangent flow. Thus, applying Lemma 3.10 finishes the proof of this case. This completes the proof of Theorem 3.16. ∎